Write down relationship between electric field and potential.
1
Expert's answer
2020-10-12T07:50:37-0400

Since the potential "\phi =\frac W q" , where "W" is energy and "q" is charge, the voltage "U=∆\phi=\frac A q" , where "A" is work, a work is equal to "A=q\times E \times ∆r" , from this follows that

"E=\frac {∆\phi}{∆r}" .